# Exploring run-and-tumble movement in confined settings through simulation

### v1

@@ 6 Processes with memory
Process memory can significantly alter the motion patterns of particles, leading to deviations from theoretical predictions like the MPLT. We are going to introduce a simple memory process in the simulations, following the next rule: the new reorientation angle of the tumble events is equal to the last one plus a constant, $\theta_{n}=\theta_{n-1}+C$. The step length distribution is exponential and the border elastic. We explore the mean and variance of the distributions for three different values of the constant $C$, as shown is the Fig. (16).
@@
The limit value of the mean path length at $l/R\rightarrow\infty$ is the Cauchy prediction, $\pi/2$, for all values of constant $C$. This occurs because the movement is essentially a billiard, and the influence of the memory process vanishes. The variance decreases because the probability distribution narrows by $l=2R$, as can be seen in Fig. (17). In this type of memory process, the distribution has two peaks, one of which decreases and the other increases as $l/R$ increases. This shows that memory processes introduce different movement patterns that deviate from the expected behavior under Markovian assumptions.
@@
Regarding the radial distribution, this presents a non-trivial form, as appreciated in Fig. (18). This shows how the reorientation mechanism of an active particle can produce distinctive distributions, hence bringing information about its underlaying laws.
Understanding process memory is not only theoretically intriguing but also practically significant. For instance, [14] demonstrated invariance properties of bacterial random walks in complex structures, suggesting that real-world bacterial motion often exhibits memory effects that need to be accounted for in predictive models. Additionally, the impact of memory on narrow escape times and sorting of active particles was explored in [20], emphasizing the practical applications in microfluidic device design and biological systems .
By integrating these insights, it becomes clear that memory effects play a fundamental role in the dynamics of confined run-and-tumble particles. This study’s exploration of simple memory processes provides a foundational understanding that can be expanded with more complex, real-world inspired mechanisms in future research. Such efforts will enhance our ability to predict and control particle behavior in various scientific and technological applications.
7 Conclusions
In this paper, we explored the dynamics of run-and-tumble motion within confined two-dimensional domains, specifically examining how various distributions and boundary conditions affect the validity of the Mean Path Length Theorem (MPLT) and the statistics of the dynamics. Through a combination of theoretical analysis and numerical simulations, we demonstrated that the MPLT holds true across a wide range of conditions, provided that the angular distribution of particle tumbling remains uniformly random. This establishes the significance of understanding motion within bounded domains across physics, ecology, and biology, and highlights the foundational role of the Mean Path Length Theorem (MPLT) in predicting and controlling particle or agent behavior in diverse environments.
We used a simulation of a 2D circular domain to model particle movement, adhering to real-world experimental conditions. By employing a constant speed for particles and exploring various step length distributions, we observe the transition of movement from diffusive to quasi-ballistic. Then, we presented an analysis on the MPLT’s applicability under diverse conditions, revealing that the theorem holds true when angular distributions are uniformly random. In fact, when particles exhibit a strong directional bias due to fixed or Gaussian-distributed angles, they tend to collide with the same region of the boundary, introducing anisotropy and thereby violating the MPLT’s foundational assumptions.
@@ 7 Conclusions
The analysis of radial distributions under different boundary conditions reveals a uniform density across all cases of $l/R$ when conditions are elastic, demonstrating a consistent linear radial distribution. The contrast with random angle boundary conditions, where the radial distribution deviates, highlights the significant impact of boundary interactions on particle distribution. This can be relevant to other physically important problems of motion in bounded systems, such as the problem of narrow escape [42], since increased spatial density near the edge increases the probability of hitting the target. Random reflections can also be important when considering reversible sticking to the boundary (see e.g. [43, 44, 45]), because if a particle sticks to the boundary, it is natural to consider the outgoing angle of release as random and uncorrelated with the incoming direction. We show in this work that the study of the statistics of movement in confined settings can bring some light to explain the mobility mechanism of some active agents, like microbia. The way in which the active particle interacts with the boundary and the reorientation mechanism produce distinctive patterns of movement and distributions, giving us information about the underlaying laws governing its movement.
@@
This work not only affirms the robustness of the MPLT under a broad spectrum of conditions but also highlights the importance of angular distribution and boundary interactions in governing the dynamics of confined motion. The implications of our research extend beyond theoretical physics, offering insights into biological processes such as bacterial chemotaxis, and the design of microfluidic devices for sorting and analyzing microscopic particles.
8 Appendix
We present here an alternative procedure for calculating the radial distribution based on [46]. Consider the case of random angle boundary condition. The distribution can be calculated as follows. Consider a part of the trajectory inside a circle or radius $R$, as showed in Fig. (19). The particle is situated at the point $x$, at distance $r$ from the centre, and moves with direction $\overrightarrow{v}$, where $|\overrightarrow{v}|=1$. In this situation, the particle will follow the trajectory $\tau$, and will collide with the boundary at the point $y$. $v_{y}$ is the normal vector at the point $y$, as showed in Fig. (19).
Figure 19: Geometry of a single collision with the boundary.
Following [46], the probability of the particle of being in a given point with a certain direction is:
@@ 8 Appendix
$\overrightarrow{v}\cdot\overrightarrow{v_{y}}=|\overrightarrow{v}|.|\overrightarrow{v_{y}}|.cos\phi,$ (26)
@@
$\frac{r}{sin\phi}=\frac{R}{sin\theta},$ (27)
@@
$|cos\phi|=\sqrt{1-sin^{2}\phi}=\sqrt{1-\frac{r^{2}}{R^{2}}sin^{2}\theta}.$ (28)
@@
$\Pi(x,\theta)=\frac{\alpha}{\sqrt{1-\frac{r^{2}}{R^{2}}sin^{2}\theta}}.$ (29)
@@
$\rho(x)=\int_{0}^{2\pi}\frac{\alpha d\theta}{\sqrt{1-\frac{r^{2}}{R^{2}}sin^{2}\theta}}=4\alpha\int_{0}^{\pi/2}\frac{\alpha d\theta}{\sqrt{1-\frac{r^{2}}{R^{2}}sin^{2}\theta}}=4\alpha K\left(\frac{r}{R}\right),$ (30)
@@
$p(r)=2\pi r\cdot 4\alpha K\left(\frac{r}{R}\right).$ (32)
After normalization, we find that, $\alpha=\frac{1}{8\pi R^{2}}$, so
@@
the same results as in eq. (23). A similar calculation can be performed for the elastic case.
Acknowledgment
R.A. acknowledges an affiliation with INDAM.

### v2

@@ VI Processes with memory
Process memory can significantly alter the motion patterns of particles, leading to deviations from theoretical predictions like the MPLT. We will introduce a simple memory process in the simulations, following this rule: the new reorientation angle of the tumble events is equal to the last one plus a constant, $\theta_{n}=\theta_{n-1}+C$. The step length distribution is exponential, and the boundary is elastic. We explore the mean and variance of the distributions for three different values of the constant $C$, as shown in Fig. (15).
@@
The limit value of the mean path length as $l/R\rightarrow\infty$ is the Cauchy prediction, $\pi/2$, for all values of constant $C$. This occurs because the movement is essentially a billiard, and the influence of the memory process vanishes. The variance decreases because the probability distribution narrows around $l=2R$, as can be seen in Fig. (16). In this type of memory process, the distribution has two peaks, one of which decreases while the other increases as $l/R$ increases. This demonstrates that memory processes introduce different movement patterns that deviate from the expected behavior under Markovian assumptions.
The memory process we chose is inspired by certain types of movement observed in bacteria, for example, the chiral movement seen in bacteria swimming close to a surface, as well as the behavior of some L-shaped active colloids. We selected this specific type of process primarily for its simplicity, which allows us to explore the fundamental aspects of memory in particle dynamics within a confined domain. However, we would like to clarify that this is not intended as a general result for all types of memory processes.
@@
Regarding the radial distribution, it presents a non-trivial form, as shown in Fig. (17). This demonstrates how the reorientation mechanism of an active particle can produce distinctive distributions, thereby providing information about its underlying laws.
Understanding process memory is not only theoretically intriguing but also practically significant. For instance, Frangipane et al. demonstrated invariance properties of bacterial random walks in complex structures, suggesting that real-world bacterial motion often exhibits memory effects that need to be accounted for in predictive models Frangipane et al. (2019). Additionally, the impact of memory on narrow escape times and the sorting of active particles was explored by Paoluzzi et al., emphasizing the practical applications in microfluidic device design and biological systems Paoluzzi, Angelani, and Puglisi (2020).
By integrating these insights, it becomes clear that memory effects play a fundamental role in the dynamics of confined run-and-tumble particles. This study’s exploration of simple memory processes provides a foundational understanding that can be expanded with more complex, real-world-inspired mechanisms in future research. Such efforts will enhance our ability to predict and control particle behavior in various scientific and technological applications.
VII Conclusions
In this paper, we explored the dynamics of run-and-tumble motion within confined two-dimensional domains, specifically examining how various distributions and boundary conditions affect the validity of the Mean Path Length Theorem (MPLT) and the statistics of the dynamics. Through a combination of theoretical analysis and numerical simulations, we demonstrated that the MPLT holds true across a wide range of conditions, provided that the angular distribution of particle tumbling remains uniformly random. This underscores the significance of understanding motion within bounded domains across physics, ecology, and biology, and highlights the foundational role of the MPLT in predicting and controlling particle or agent behavior in diverse environments.
We used a simulation of a 2D circular domain to model particle movement, adhering to real-world experimental conditions. By employing a constant speed for particles and exploring various step length distributions, we observed the transition of movement from diffusive to quasi-ballistic. We then presented an analysis of the MPLT’s applicability under diverse conditions, revealing that the theorem holds true when angular distributions are uniformly random. In fact, when particles exhibit a strong directional bias due to fixed or Gaussian-distributed angles, they tend to collide with the same region of the boundary, introducing anisotropy and thereby violating the MPLT’s foundational assumptions.
@@ VII Conclusions
The analysis of radial distributions under different boundary conditions reveals a uniform density across all cases of $l/R$ when conditions are elastic, demonstrating a consistent linear radial distribution. The contrast with random angle boundary conditions, where the radial distribution deviates, highlights the significant impact of boundary interactions on particle distribution. This is relevant to other physically important problems of motion in bounded systems, such as the problem of narrow escape Holcman and Schuss (2016), since increased spatial density near the edge raises the probability of hitting the target. Random reflections can also be important when considering reversible sticking to the boundary (see, e.g., Grebenkov (2017); Grebenkov and Kumar (2022); Grebenkov (2023)), because if a particle sticks to the boundary, it is natural to consider the outgoing angle of release as random and uncorrelated with the incoming direction. We show in this work that studying the statistics of movement in confined settings can help explain the mobility mechanisms of some active agents, like microbes. The way in which the active particle interacts with the boundary and the reorientation mechanism produces distinctive patterns of movement and distributions, giving us information about the underlying laws governing its movement.
@@
This work not only affirms the robustness of the MPLT under a broad spectrum of conditions but also highlights the importance of angular distribution and boundary interactions in governing the dynamics of confined motion. The implications of our research extend beyond theoretical physics, offering insights into biological processes such as bacterial chemotaxis and the design of microfluidic devices for sorting and analyzing microscopic particles.
Acknowledgements.
R.A. acknowledges an affiliation with INDAM.
Data Availability Statement
The data that support the findings of this study are available within the article.
Author contribution statement
D.J.Z.: Visualization, Validation, Investigation, Formal analysis, Data curation, Writing - original draft. R.A.: Writing - review & editing, Supervision, Resources, Methodology, Funding acquisition, Conceptualization.
Declaration of competing interest
The authors have no conflicts to disclose.
Appendix A Alternative calculation of distributions.
We present here an alternative procedure for calculating the radial distribution based on Evans (2001). Consider the case of random angle boundary conditions. The distribution can be calculated as follows: consider a portion of the trajectory inside a circle of radius $R$, as shown in Fig. (18). The particle is situated at the point $x$, at distance $r$ from the center, and moves in the direction $\overrightarrow{v}$, where $|\overrightarrow{v}|=1$. In this situation, the particle will follow the trajectory $\tau$, and will collide with the boundary at the point $y$. $v_{y}$ is the normal vector at the point $y$, as shown in Fig. (18).
Following Evans (2001), the probability of the particle of being in a given point with a certain direction for the case of random boundary conditions is:
@@ Acknowledgements.
$\overrightarrow{v}\cdot\overrightarrow{v_{y}}=|\overrightarrow{v}|\times|\overrightarrow{v_{y}}|\times\cos\phi,$ (34)
@@
$\frac{r}{\sin\phi}=\frac{R}{\sin\theta},$ (35)
@@
$|\cos\phi|=\sqrt{1-\sin^{2}\phi}=\sqrt{1-\frac{r^{2}}{R^{2}}\sin^{2}\theta}.$ (36)
@@
$\Pi(x,\theta)=\frac{\alpha}{\sqrt{1-\frac{r^{2}}{R^{2}}\sin^{2}\theta}}.$ (37)
Figure 18: Geometry of a single collision with the boundary.
@@
$\begin{split}\rho(x)&=\int_{0}^{2\pi}\frac{\alpha d\theta}{\sqrt{1-\frac{r^{2}}{R^{2}}\sin^{2}\theta}}=4\alpha\int_{0}^{\pi/2}\frac{\alpha d\theta}{\sqrt{1-\frac{r^{2}}{R^{2}}\sin^{2}\theta}}\\ &=4\alpha K\left(\frac{r}{R}\right),\end{split}$ (38)
@@
$p(r)=2\pi r\times 4\alpha K\left(\frac{r}{R}\right).$ (40)
After normalization, we find that, $\alpha=1/8\pi R^{2}$, so
@@
the same results as in eq. (31).
A similar calculation can be performed for the elastic case. For elastic collisions $\Pi(x,\theta)=\alpha$, since the invariant measure is uniform, and the computation is simple.
[…]
The radial probability is
$p(r)=2\pi r\times 2\pi\alpha=4\pi\alpha r,$ (43)
which after normalization becomes
$p(r)=\frac{2r}{R^{2}},$ (44)
the same result as in eq. (27).